Average total cost is
a. AFC + (TVC/Q)
b. TC/Q   
c. (TFC/Q) + (TVC/Q)
d. AFC + AVC
e. all of the above

Please Help!
Thank you!
Answer #1

Answer : The answer is option e.

AFC = TFC / Q

AVC = TVC / Q

We can find out ATC in two ways. One is :

Average Total Cost (ATC) = AFC + AVC.

And another way is :

Average Total Cost = TC / Q

As here options a, b, c and d all are correct, hence the answer is option e.